# It’s Not Easy Being Green: On the Energy Efficiency of Programming Languages

### v2

@@ 4.4.1. Number of Active Cores
4.4.2. Memory Activity
Previous research has shown that roughly 40% of dynamic random access memory (DRAM) energy consumption is constant and required to refresh memory cells to keep the system running properly, while the remaining 60% is related to read and write activity (Seol et al., 2016). Section 3.2.2 introduces last level cache (LLC) misses as a better proxy for memory activity: any read or write request that cannot be fulfilled by the LLC will have to go to DRAM. Figure 12(b) shows a strong linear correlation between LLC misses and increased power draw $\left(R^{2}=0.93\right)$.
@@ 4.4.2. Memory Activity
4.5. Application Implementations
Figure 13. The final model incorporating application implementations (§4.5). Application implementations notably dictate the number of cores used by the program and the use of third-party libraries. These implementation choices are not properties of the programming language itself.
[…]
Our previous working causal diagram, adding “Application Implementation”.
Finally, Figure 13 adds the application implementations themselves, which have the most direct impact on energy consumption. Implementations dictate the level of parallelism, the algorithms used, and the data structures employed. We discuss two factors, concurrency and third-party library usage, below.
4.5.1. Concurrency
Section 4.4.1 asserts that the number of cores used by an application is a major factor in energy consumption. Therefore, comparing benchmark implementations in different languages that use different numbers of cores is not a fair comparison. Level of parallelism imbalances are the main reason for the reported performance discrepancies between JavaScript and TypeScript. The TypeScript version of the mandelbrot benchmark takes 21×21$\times$21 × longer to run than its JavaScript implementation. Its execution is fully sequential, while the JavaScript version uses 28 cores on average.
@@ 4.5.1. Concurrency
4.5.2. Third-Party Libraries
@@
Using third-party libraries also allows for easily crossing language boundaries, for instance to access lower-level system utilities or to obtain greater performance. In Python, using NumPy can yield up to 60,000×60,000$\times$60 , 000 × faster code when multiplying matrices (Leiserson et al., 2020; Berger et al., 2023). On a small micro-benchmark multiplying $000\times 10,$ 000 × 10 , 000 matrics, we find that NumPy is 84×84$\times$84 × faster than a naive three-loop C++ implementation, with $\times4$ lower energy consumption. Of course, this is because NumPy’s underlying matrix-multiply implementation is written in a low-level language and leverages many optimizations such as parallelism, vector instructions, blocking, or hardware-specific instructions. To some degree, using NumPy and other third-party libraries has become the “correct” way of writing Python code: leveraging low-level languages for compute-intensive tasks, while keeping Python’s simplicity. Even C++ may cross over to code originally written using C, Fortran, or direct Assembly for performance when using BLAS.
4.6. Other Factors
@@
There are many other factors that may affect energy consumption. CPU voltage and frequency scaling has a direct impact on power consumption, and there is active research focusing on reducing or increasing frequency based on memory activity or other factors to draw less power for a comparatively smaller performance penalty (Weissel and Bellosa, 2002). Specific processor model and architecture also have an impact on power draw, with newer processors generally being more energy-efficient. These factors are unrelated to choice of programming languages and are controlled for in our experiments.
5. Related Work
@@
Previous sections, notably Section 2, discuss the main line of research relating programming languages and energy consumption (Couto et al., 2017; Pereira et al., 2017, 2021). These studies discuss and rank programming languages based on execution time, energy consumption, and memory usage. Contrary to Pereira et al.’s results, we show in this paper that the interplay between these properties is complex and depends on many factors not taken into account in their work. Gordillo et al. (Gordillo et al., 2024) and Koedijk et al. (Koedijk and Oprescu, 2022) build on this line of research using direct hardware power measurements but suffer from similar limitations.
@@
The processor’s operating voltage and frequency also have an impact on energy consumption. This impact has been studied in the context of trading off speed for reduced energy consumption, where compiling for speed yields better energy efficiency in the general case (Yuki and Rajopadhye, 2013). For some specific applications, such as sparse matrix computations, compile-time techniques may decrease energy usage without impact to execution time by leveraging load imbalance (Chen et al., 2005). Other work has shown that using characteristics of a program during execution to reduce the processor’s voltage or frequency can yield higher energy efficiency for a comparably smaller performance degradation (Hsu and Feng, 2005). For example, performance counter information can be used to scale the processor’s frequency (Weissel and Bellosa, 2002). These approaches are orthogonal to the choice of programming language and could be applied in many programming environments.
@@
Graphics processing units (GPUs) are becoming increasingly prevalent in modern computing systems, notably in artificial intelligence workloads. Several studies have examined the energy efficiency of GPUs from a software perspective, notably accounting for the energy expenditure of tensors (Babakol and Liu, 2024) and comparing the energy consumption of different machine learning frameworks (Georgiou et al., 2022). Extending our work to incorporate a model for energy consumption of GPUs is a potential avenue for future work.
@@
Previous research also addresses the energy consumption of different data structures in certain languages, notably Java collections (Hasan et al., 2016; Pereira et al., 2016; Oliveira et al., 2021), or Python data manipulation libraries (Nahrstedt et al., 2024). Data structures can be interesting candidates to study with an energy-focused lens as they can have a substantial impact on locality and memory activity, which are factors in energy consumption (as Section 4.4.2 describes). In many cases, however, programmers only have a limited choice of data structures that would be a good fit for their use case or algorithm.
@@
Finally, past work has also investigated energy efficiency of concurrent programs, focusing on Haskell’s data sharing primitives (Lima et al., 2016) or Java’s thread management constructs (Pinto et al., 2014), or analyzing the power to performance tradeoffs in lock-based synchronization (Falsafi et al., 2016). As Section 4.4.1 argues, these studies confirm that the power usage patterns of the processor and the machine as a whole are more complex with parallel execution.
6. Conclusion
@@
This paper presents a detailed causal model exploring the complex relationship between choice of programming language and energy consumption. It shows some of the many factors at play, notably distinguishing implementation from programming language and establishing the number of active cores and memory activity as two important factors in power draw.
@@
Using this causal model, we investigate and explain anomalies in previous research, finding that many factors such as parallelism level, benchmark implementation specifics, or language implementation properties must be taken into account for a fair comparison. Our results suggest that the choice of programming language has no significant impact on energy consumption beyond execution time. Programmers aiming to reduce energy consumption can do so even in “inefficient” programming languages like Python by using faster language implementations, employing faster and more parallel algorithms, and using fast third-party (native) libraries.
@@
The software and benchmarks that support this paper’s experiments are available on GitHub at github.com/nicovank/Energy-Languages.
Acknowledgements.
We thank Prashant Shenoy, Walid Hanafy, Noman Bashir, Mehmet Savasci, and Abel Souza for their help getting started with energy measurements using RAPL, their continuous feedback, and for providing access to an energy-measurement-ready machine. We thank Eunice Jun for suggesting exploration of causal models in this work. Emery Berger holds concurrent appointments as a Professor of Computer Science at the University of Massachusetts Amherst and as an Amazon Scholar. This paper describes work performed at the University of Massachusetts Amherst and is not associated with Amazon.

### v3

@@ IV-D1 Number of Active Cores
IV-D2 Memory Activity
Previous research has shown that roughly 40% of dynamic random access memory (DRAM) energy consumption is constant and required to refresh memory cells to keep the system running properly, while the remaining 60% is related to read and write activity [35]. Section III-B2 introduces last level cache (LLC) misses as a better proxy for memory activity: any read or write request that cannot be fulfilled by the LLC will have to go to DRAM. Figure 9(b) shows a strong linear correlation between LLC misses and increased power draw $\left(R^{2}=0.93\right)$.
@@ IV-D2 Memory Activity
IV-E Application Implementations
Finally, we add the application implementations themselves, which have the most direct impact on energy consumption. Figure 1 presents the final causal model developed in this paper. Application implementation decisions dictate the level of parallelism, the algorithms used, and the data structures employed. We discuss two factors, concurrency and third-party library usage, below.
IV-E1 Concurrency
Section IV-D1 asserts that the number of cores used by an application is a major factor in energy consumption. Therefore, comparing benchmark implementations in different languages that use different numbers of cores is not a fair comparison. Level of parallelism imbalances are the main reason for the reported performance discrepancies between JavaScript and TypeScript. The TypeScript version of the mandelbrot benchmark takes 21×21$\times$21 × longer to run than its JavaScript implementation. Its execution is fully sequential, while the JavaScript version uses 28 cores on average.
@@ IV-E1 Concurrency
IV-E2 Third-Party Libraries
@@
Using third-party libraries also allows for easily crossing language boundaries, for instance to access lower-level system utilities or to obtain greater performance. In Python, using NumPy can yield up to 60,000×60,000$\times$60 , 000 × faster code when multiplying matrices [36, 22]. On a small micro-benchmark multiplying $000\times 10,$ 000 × 10 , 000 matrics, we find that NumPy is 84×84$\times$84 × faster than a naive three-loop C++ implementation, with $\times4$ lower energy consumption. Of course, this is because NumPy’s underlying matrix-multiply implementation is written in a low-level language and leverages many optimizations such as parallelism, vector instructions, blocking, or hardware-specific instructions. To some degree, using NumPy and other third-party libraries has become the “correct” way of writing Python code: leveraging low-level languages for compute-intensive tasks, while keeping Python’s simplicity. Even C++ may cross over to code originally written using C, Fortran, or direct Assembly for performance when using BLAS.
IV-F Other Factors
@@
There are many other factors that may affect energy consumption. CPU voltage and frequency scaling has a direct impact on power consumption, and there is active research focusing on reducing or increasing frequency based on memory activity or other factors to draw less power for a comparatively smaller performance penalty [37]. Specific processor model and architecture also have an impact on power draw, with newer processors generally being more energy-efficient. These factors are unrelated to choice of programming languages and are controlled for in our experiments.
V Threats to Validity
This study largely focuses on reproducing Pereira et al.’s experiments, which rely on a specific set of benchmarks. We mitigate this threat to external validity by not only focusing on general run-time properties of those programs rather than their absolute execution time, but also by verifying results across two additional benchmark suites.
Experiments in this paper are conducted on a server architecture, maintaining comparability with prior work. While we expect our conclusions to generalize to other architectures, results may differ on mobile, embedded systems, or architectures mixing energy-saving and high-performance cores.
VI Related Work
@@
Previous sections, notably Section II, discuss and critique the main line of research relating programming languages and energy consumption [9, 10, 11]. These studies discuss and rank programming languages based on execution time, energy consumption, and memory usage. Follow-on studies [38, 39] use external energy measurements rather than the on-chip measurements used here and in Pereira et al., but still suffer from the same limitations as the original work.
@@
Processor operating voltage and frequency have an impact on energy consumption. This impact has been studied in the context of trading off speed for reduced energy consumption, where compiling for speed yields better energy efficiency in the general case [40]. For specific applications, such as sparse matrix computations, compile-time techniques may decrease energy usage without impacting execution time by leveraging load imbalance [41]. Other work has shown that using characteristics of a program during execution to reduce the processor’s voltage or frequency can yield higher energy efficiency for a comparably smaller performance degradation [42]. For example, performance counter information can be used to scale the processor’s frequency [37]. These approaches are orthogonal to the choice of programming language and could be applied in many programming environments.
@@
Graphics processing units (GPUs) are increasingly prevalent in modern computing systems, notably in artificial intelligence workloads. Past work has examined the energy efficiency of GPUs from a software perspective, notably accounting for the energy expenditure of tensors [43] and comparing the energy consumption of different machine learning frameworks [44]. Extending our work to incorporate a model for energy consumption of GPUs is a potential avenue for future work.
@@
Previous research also addresses the energy consumption of different data structures in certain languages, notably Java collections [45, 46, 47], or Python data manipulation libraries [48]. Data structures can be interesting candidates to study through an energy-focused lens as they can have a substantial impact on locality and memory activity, which, as Section IV-D2 describes, are factors in energy consumption. However, programmers often have a limited choice of data structures that fit for their use case or algorithm.
@@
Finally, past work has also investigated energy efficiency of concurrent programs, focusing on Haskell’s data sharing primitives [49] or Java’s thread management constructs [50], or analyzing the power to performance tradeoffs in lock-based synchronization [51]. As Section IV-D1 argues, these studies confirm that the power usage patterns of the processor and the machine as a whole are more complex with parallel execution.
VII Conclusion
@@
This paper presents a detailed causal model exploring the complex relationship between choice of programming language and energy consumption. It captures some of the factors at play, notably distinguishing implementation from programming language, and establishing the number of active cores and memory activity as important factors in power draw.
@@
Using this causal model, we investigate and explain anomalies in previous research, finding that many factors such as parallelism level, benchmark implementation specifics, or language implementation properties must be taken into account for a fair comparison. Our results suggest that the choice of programming language has no significant impact on energy consumption beyond execution time. Programmers aiming to reduce energy consumption can do so by focusing on performance optimizations. This strategy is possible even in “inefficient” programming languages like Python by using faster language implementations, employing faster and more parallel algorithms, and using native libraries.
@@
The software, benchmarks, and data that support this paper’s experiments are available on GitHub at github.com/nicovank/Energy-Languages.
Acknowledgments
We thank Prashant Shenoy, Walid Hanafy, Noman Bashir, Mehmet Savasci, and Abel Souza for their help getting started with energy measurements using RAPL, their continuous feedback, and for providing access to an energy-measurement-ready machine. We thank Eunice Jun for suggesting exploration of causal models in this work.
Emery Berger holds concurrent appointments as a Professor of Computer Science at the University of Massachusetts Amherst and as an Amazon Scholar. This paper describes work performed at the University of Massachusetts Amherst and is not associated with Amazon.